The power of the heat fan is 2.00 kW. How much coal should be burned per hour to produce the corresponding heat when the calorific value of the coal is 28.4 MJ/kg?


1
Expert's answer
2021-10-11T16:45:10-0400

Power of 1W is equal to 1J in 1s. Thus, in one hour (3600 seconds) the coal should suplement the following amount of energy:


"Q = 2\times 10^3\dfrac{J}{s}\cdot 3600s = 7.2\times 10^6J = 7.2MJ"

The mass of coal is given as follows:


"m = \dfrac{Q}{q}"

where "q = 28.4MJ\/kg" is the calorific value of the coal. Thus, obtain:


"m = \dfrac{7.2MJ}{28.4MJ/kg} \approx 0.253kg"

Answer. 0.253 kg.